OpenStudy (anonymous):

Given the equation P2 = A3, what is the orbital period, in years, for the planet Jupiter? (Jupiter is located 5.2 AU from the sun.) 

OpenStudy (jfraser):

plug and solve\[P^2 = A^3\]so\[P = \sqrt{A^3} = \sqrt{(5.2AU)^3}\]

OpenStudy (anonymous):

I got 11.83 so the choices i have would be 11 the other options are 5 years 11 years 110 years 500 years @JFraser

OpenStudy (jfraser):

take the cube of 5.2 and take the square root of that. the answer is one of those